# DS-VTON: High-Quality Virtual Try-on via Disentangled Dual-Scale Generation

### v1

@@ B.1 Ablation on removing cross-attention layers in U-Net
Appendix C Discussions
@@
In addition to our proposed refinement approach, we also explored an alternative refinement strategy inspired by SDXL [Podell et al., 2025]. In SDXL, the refinement process involves training a separate model that focuses specifically on denoising the final 200 steps of the diffusion process. During inference, noise corresponding to timestep 200 is added to the output of the first-stage model, and the refinement model is then used to denoise and generate the final result. We experimented with two variations of this approach. In the first variant, the first-stage and refinement-stage outputs share the same resolution. However, as shown in Subsection 4.3, directly applying a mask-free strategy at high resolution inevitably leads to structural distortions, which cannot be effectively corrected by the refinement stage. This led us to the second variant, which we refer to as the noising-denoising strategy. Here, a low-resolution result is first generated and then upsampled to high resolution. Noise is added at this stage (corresponding to a specific timestep), and the refinement model performs denoising from that point onward. While this method can sometimes produce visually plausible results when the target patterns are simple, as shown in Table 5 and Figure 7, its overall performance remains suboptimal. The reason, we believe, lies in the difference between data distributions involved. Although our residual-guided denoising strategy also begins with adding noise to the low-resolution output, it differs fundamentally in how it relates the two stages. In the noising-denoising approach, the refinement model learns to denoise samples drawn solely from the high-resolution distribution, and it lacks an explicit mechanism to relate this with the low-resolution inputs. In contrast, our residual-guided denoising bridges the gap between the low-resolution and high-resolution distributions. This connection is key: it allows the model to better capture the transformation between the two distributions involved. Importantly, while the original diffusion process constructs a mapping from a simple, tractable distribution (such as Gaussian noise) to a complex data distribution, the residual-guided refinement builds a direct bridge between two complex distributions.
Figure 7: Comparison between our baseline and the noising-denoising variant.
Table 5: Comparison of different refinement strategies on the VITON-HD dataset.
Version FID $\downarrow$ KID $\downarrow$
Noising-denoising way 8.98 0.65
DS-VTON 8.24 0.31
@@
As discussed in Subsection A.1, one key limitation of our method lies in the data generation process. Due to the reliance on synthesized person images, the model may inadvertently learn to alter regions unrelated to the garment (e.g., hair, accessories, or background). While this issue is not severe in most cases, we acknowledge it as a limitation and consider improving data disentanglement and identity preservation an important direction for future work. Another limitation stems from the use of fixed coefficients $\alpha$ and $\beta$ to initialize the high-resolution refinement stage. Although this static strategy proves effective, it may be overly rigid. In future work, we plan to investigate adaptive or learnable coefficient scheduling mechanisms, which could offer more flexible and content-aware refinement during generation.
Appendix D More experiment results
In this section, we present additional qualitative comparison results on the DressCode [Morelli et al., 2022] dataset and more comparison results on the VITON-HD [Choi et al., 2021] dataset.

### v2

@@ A.2.1 Ablation on removing cross-attention layers in U-Net
Table 4: Quantitative comparisons on the VITON-HD dataset under the paired setting.
Method FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$
OOTDiffusion 6.47 1.24 0.88 0.08
IDM-VTON 5.84 0.77 0.87 0.06
CatVTON 5.70 0.50 0.88 0.09
Leffa 5.76 0.55 0.89 0.06
FitDiT 7.27 0.73 0.84 0.09
DS-VTON (direct) 4.75 0.43 0.90 0.05
DS-VTON (train-way) 5.23 0.31 0.89 0.06
@@
A.3 Discussions
Figure 7: Comparison of qualitative results between our baseline and the SDXL variant. Quantitatively, the SDXL variant achieves an FID/KID of 8.98/0.65, while our DS-VTON achieves 8.24/0.31.
@@ A.3 Discussions
We refer to the refinement method in SDXL [Podell et al., 2025] as the SDXL strategy. During training, a separate model is trained to denoise only the final 200 steps of the high-resolution diffusion process. At inference, a low-resolution result is generated, upsampled, and injected with Gaussian noise at timestep 200, after which the refinement model denoises to produce the final output. Although this method can yield visually plausible results for simple patterns, its overall performance is inferior, as shown in Figure 7. The reason, we believe, lies in the difference between data distributions involved. Although our blend-refine diffusion process also begins with adding noise, it differs fundamentally in how it relates the two stages. In SDXL, the refinement model learns to denoise samples drawn solely from the high-resolution distribution, and it lacks an explicit mechanism to relate this with the low-resolution result. In contrast, our blend-refine diffusion bridges the gap between the low-resolution and high-resolution distributions. This connection is key: it allows the model to better capture the transformation between the two distributions involved. Importantly, while the original diffusion process constructs a mapping from a simple, tractable distribution (such as Gaussian noise) to a complex data distribution, the blend-refine diffusion builds a direct bridge between two complex distributions.
Difference between Our Blend-Refine Process and the DCI-VTON Refinement Branch.
DCI-VTON also adds noise to the first-stage result and attempts to recover the final output in its second-stage refinement branch. However, it does not explicitly transition the data distribution from the first stage to the second. Its second stage contains two branches: a refinement branch and a reconstruction branch. In the refinement branch, when the added noise is weak, the noisy input is nearly identical to the first-stage result, so the denoised output remains overly similar to the input and struggles to reach the desired final distribution. This is why DCI-VTON requires two branches and adopts a perceptual (VGG) loss rather than a pixel-wise loss in the refinement branch. The refinement branch provides only implicit guidance, while the reconstruction branch handles explicit Gaussian noise prediction. By contrast, DS-VTON needs only a single branch. Our blend-refine denoising explicitly transfers the distribution from coarse to fine by altering the noise formulation itself, rather than relying on pure Gaussian noise. This change enables a principled transition between the low- and high-resolution distributions and provides stronger controllability—an essential design choice of our dual-scale framework. To illustrate, suppose we follow DCI-VTON’s approach. Let $\tilde{\mathbf{x}}_{r}$ denote the low-resolution output, $\bm{\epsilon}$ the noise, and $\mathbf{x}_{r}$ the high-resolution ground truth. From a data distribution perspective, if we adopt pure Gaussian noise as the only perturbation mechanism, there is a fundamental mismatch between the training and inference processes. During training, at any timestep $t$, the noisy latent is constructed by adding noise to $\tilde{\mathbf{x}}_{r}$, i.e., it is solely a function of $\tilde{\mathbf{x}}_{r}$ and $\bm{\epsilon}$. However, this setup does not align with inference. In inference, regardless of how the latent is initialized (whether from pure Gaussian noise or some other distribution), the moment a single denoising step is performed, the latent becomes correlated with the target distribution $\mathbf{x}_{r}$. This stands in stark contrast to training, where the noisy latent at each timestep is always constructed based on $\tilde{\mathbf{x}}_{r}$. The key difference is that we modify the noise formulation: rather than using pure Gaussian noise, our blend-refine denoising bridges the low- and high-resolution data distributions, providing strong controllability and serving as a fundamental design choice in our dual-scale framework.
@@
As discussed in Subsection A.1.1, one key limitation of our method lies in the data generation process. Due to the reliance on synthesized person images, the model may inadvertently learn to alter regions unrelated to the garment (e.g., hair, accessories, or background). While this issue is not severe in most cases, we acknowledge it as a limitation and consider improving data disentanglement and identity preservation an important direction for future work. Another limitation stems from the use of fixed coefficients $\alpha$ and $\beta$ to initialize the high-resolution refinement stage. Although this static strategy proves effective, it may be overly rigid. In future work, we plan to investigate adaptive or learnable coefficient scheduling mechanisms, which could offer more flexible and content-aware refinement during generation.
A.4 The Use of Large Language Models
The Large Language Model was used solely for refining the text and improving the clarity and readability of the paper. It did not contribute to the research ideation or experimental design.
A.5 More experiment results
In this section, we present additional qualitative comparisons on the DressCode [Morelli et al., 2022] dataset, more results on the VITON-HD [Choi et al., 2021] dataset, and additional in-the-wild examples.
Figure 8: More results on in-the-wild scenarios.